OpenStudy (anonymous):

12x^2+6x-8. Vertex is -2, which is minimum value.

OpenStudy (anonymous):

Vertex is -2?

OpenStudy (anonymous):

x coordinate of vertex: -b/2a = -6/24 = -1/4

OpenStudy (anonymous):

substitute -1/4 12(-1/4)^2+6(-1/4)-8 = -8.75

OpenStudy (anonymous):

vertex is (-0.25, -8.75) Minimum value is -8.75